A man uses 1/8 of his salary after tax on house rent and a 3/5 on transport, food and other household items. He then reserves 5/8 of the remainder on leisure and incidentals while saving the rest. If he saves $37135 every monthHow much is his Salary after tax ? (round to the nearest whole number )


1
Expert's answer
2021-03-04T23:47:14-0500

Let salary be x

"1\/8*x=0.125x"

"3\/5*x=0.6x"

"X-(0.125x+0.6x)=0.275x"

Balance"=0.275x"

"5\/8*0.275x=0.172x"

"0.275x-0.172x=0.103x"

"0.103x=37135"

"x=360534"

$"360534"